# Triple and binary evolution pathways play equally significant roles in LISA double white dwarf formation

### v1

@@ 3.3 Ejected tertiary
About $27.0\,\%\pm 0.4\,\%$ of Galactic double white dwarfs from triples lost the tertiary star. The unbinding of the tertiary companion primarily occurs due to the following reasons. First, during a CE phase the inner binary’s orbit shrinks and loses angular momentum, while the prompt mass loss during the CE unbinds the outer orbit. Second, if the inner binary widens due to mass transfer or stellar winds the system becomes less hierarchical and dynamically unstable. Dynamically unstable orbits lead to chaotic evolution which may eject an object from the system. We provide a specific example of this process below.
Panel (b) of Figure 6 shows an example of a system where a tertiary gets ejected after a CE in the inner binary. The system starts with $m_{1}=1.39\,\mathrm{\mathrm{M_{\odot}}}$ and $m_{2}=1.39\,\mathrm{\mathrm{M_{\odot}}}$ as the inner binary components and $m_{3}=0.31\,\mathrm{\mathrm{M_{\odot}}}$ as the tertiary star. The inner binary is initially eccentric with an inner eccentricity $\mathrm{e_{1}}=0.56$ and an inner semi-major axis $\mathrm{a_{1}}=50.1\,\mathrm{au}$. The tertiary orbit is wide, with an outer semi-major axis $\mathrm{a_{2}}\approx 1690\,\mathrm{au}$ and is eccentric with $\mathrm{e_{2}}=0.81$. The mutual inclination between the inner and outer orbits is 112. The inner binary masses are greater than the tertiary star so they evolve on a shorter timescale. The CE occurs in the inner binary during the AGB phase after about 3.6 Gyr. The CE reduces the inner binary semi-major axis to $10^{-2}\,\rm{au}$ and unbinds the outer tertiary. The remaining binary components evolve into a double white dwarf system. This double white dwarf binary emits gravitational-waves and later enters the LISA frequency bandwidth after $1\,\rm{Myr}$. In this example, even though the binary does not have a bound tertiary star by the time it enters the LISA frequency bandwidth, the tertiary plays a role in shrinking the binary before the CE episode.
Refer to captionRefer to caption
Figure 7: Schematic diagram for an example system which undergoes a TCE. In this scenario, the massive star transfers mass onto the inner binary, leading to its merger and the formation of a rejuvenated star. This rejuvenated star later enters the LISA frequency bandwidth along with the tertiary star. The legends are similar to those in Figure 2. In addition, the yellow circle represents a star in the asymptotic giant branch (AGB) phase.
@@ 3.4 Triple common envelope
About $5.5\,\%\pm 0.1\,\%$ of of Galactic double white dwarfs from triples undergo a phase of TCE evolution before entering the LISA frequency band. We identify two possible evolutionary outcomes of a TCE event in our simulations. First, the merger of the inner binary: the inner binary may merge to form a rejuvenated star. The new star forms a binary with the third star which later enters LISA frequency bandwidth. Second, ejection of one component: One of the stars may become unbound from the system, leaving behind a binary formed by the remaining two stars. This binary may enter the LISA band. We explain one of the examples below (see Hamers et al. 2022 for a more detailed study of TCE).
Figure 7 shows an example of a TCE where the inner binary merges to form a new binary. The system starts with $m_{1}=1.18\,\mathrm{\mathrm{M_{\odot}}}$, $m_{2}=0.54\,\mathrm{\mathrm{M_{\odot}}}$, and a relatively massive tertiary $m_{3}=5.59\,\mathrm{M_{\odot}}$. The inner binary is initially circular, with a semi-major axis $\mathrm{a_{1}}=0.05\,\mathrm{au}$. The tertiary orbit is compact, with an outer semi-major axis $\mathrm{a_{2}}=25\,\mathrm{au}$ and highly eccentric with $\mathrm{e_{2}}=0.94$. The mutual inclination between the inner and outer orbit is 160. Since the tertiary star is massive compared to the inner binary masses, it is the first to reach the AGB phase, in about 92 Myr. The inner binary components are still on the main-sequence.
@@
After approximately $2\,\rm{Gyr}$, the rejuvenated star evolves into a red giant which transfers mass onto the white dwarf, initiating another CE phase. The end of this CE phase leaves a circular and compact double white dwarf in a circular orbit with a semi-major axis of $\sim 9\times 10^{-3}\,\mathrm{au}$. After a few million years the binary enters the LISA frequency bandwidth due to gravitational-wave emission. Here, TCE plays a significant role as it leads to merging the inner binary. The binary with the post-merger rejuvenated star later enters the LISA frequency bandwidth. Hence, the tertiary star plays a vital role in the evolution of the triple to LISA frequency bandwidth.
@@ 3.5 Inner binary channel
The inner binary channel is similar to the isolated binary channel, where the binary evolves without a notable contribution from the tertiary star. About $39.0\,\%\pm 0.7\,\%$ of the Galactic double white dwarfs from triples in our simulations follow this channel. Here the inner binary undergoes phases of CE before forming a short-period double white dwarf. This double white dwarf later enters the LISA frequency bandwidth with either a wide tertiary or as a binary with an ejected tertiary (for details on ejected tertiary, see Section 3.3).
4 LISA double white dwarfs: isolated binary vs. triple evolution
Building on the method in Section 2.3, where we construct the Galactic population by combining MSE with a Milky Way-like galaxy from TNG50, we obtain a total of $\sim 7.2\times 10^{6}$ Galactic double white dwarfs originating from triples. To compare this result with the isolated binary channel, we adjust our simulation as described in Appendix A. We emphasize that the initial distribution of the isolated binary population is constructed to represent truly isolated binaries and is different from the initial properties sampled to assemble the inner binaries in triple systems (cf. Section 2.2). In the isolated binary case, we find that about $\sim 3.8\times 10^{6}$ double white dwarfs are produced through isolated binary evolution, yielding a total of $\sim 1.1\times 10^{7}$ Galactic double white dwarfs emitting gravitational waves within the LISA bandwidth. Thus, approximately $65\,\%$ ($\sim 7.2\times 10^{6}$) of all Galactic double white dwarf binaries is originated from triple systems as illustrated in Figure 8. We note that the isolated binary evolution channel only yields circular binaries. In contrast, the triple channel generates a small fraction ($3\times 10^{-6}$) of eccentric binaries, which we will discuss below.
Notably, we find that among all the double white dwarfs initially in triples only about half ($\sim$57%) retain a bound tertiary. In the remaining systems the third star either becomes unbound or merges with another star within the system to form a binary. Importantly, all double white dwarfs that retain their tertiary companion in our simulations have the tertiary in a relatively wide orbit. We discuss the potential for detecting the presence of the tertiary companion based on LISA data in Section 4.4.
We recall that we also generated a comparative simulation where the Galactic population only follows the isolated binary channel (i.e., entirely excluding triple channels). In this simulation we obtain a total of $\sim 9.4\times 10^{6}$ double white dwarfs emitting at LISA frequencies. All results are summarized in Table 1.
@@
Category $N(f<10^{-4}\,\textrm{Hz})\times 10^{6}$ $N(\rho>7)\times 10^{3}$
Triple $7.20\pm 0.13$ $10.9$
Binary $3.80\pm 0.10$ $6.5$
Triple + Binary $11.00\pm 0.16$ $17.4$
Only binaries ($f_{t}=0$) $9.40\pm 0.12$ $14.4$
Table 1: The table presents the estimated counts of Galactic LISA double white dwarfs from different formation channels: Triple (originating from triple systems), Binary (originating from isolated binaries), Triple + Binary (the combined population from both channels), and Only binaries (isolated binaries with no triples, i.e., a triple fraction $f_{t}=0$). Errors represent scaled up the fractional Poisson error from the intrinsic population evolved with MSE.
@@ 4.1.1 Circular systems
Double white dwarfs in the LISA band are typically millions of years away from merging. They are continuous, quasi-monochromatic gravitational-wave sources for LISA. Describing these signals requires a set of eight parameters, typically chosen as $\{{\mathcal{A}_{\rm gw},f_{\rm gw},\dot{f}_{\rm gw},\lambda,\beta,\iota,\psi,\phi_{0}}\}$ (LISA Consortium Waveform Working Group et al., 2023). Here, $\mathcal{A}_{\rm gw}$ represents the gravitational-wave amplitude, $f_{\rm gw}$ and $\dot{f}_{\rm gw}$ denote the gravitational-wave frequency and its time derivative (or chirp), $(\lambda,\beta)$ correspond to the ecliptic longitude and latitude, respectively, $\iota$ is the (inner) binary inclination angle with respect to the line-of-sight, $\psi$ is the gravitational-wave polarization angle, and $\phi_{0}$ represents the binary’s initial phase. Our population synthesis models provide binary parameters such as component masses, orbital periods and eccentricities, sky positions and distances. We use these to derive gravitation-wave parameters as follows. As discussed above, the overwhelming majority of double white dwarfs in our simulations are circularized.
The gravitational-wave frequency of a circular binary is twice its orbital frequency $f_{\rm orb}$
@@
where $G$ and $c$ are the gravitational constant and speed of light respectively. The amplitude is set by the source’s distance $d$ and chirp mass
@@
Equations (3), (4), and (6) define the first three parameters of the set. The ecliptic coordinates $(\lambda,\beta)$ are inherited from the TNG50 Milky Way-like galaxy where the binary was seeded, while the remaining three parameters are assigned randomly: $\iota$ is sampled from a uniform distribution in $\cos\iota$, and $\psi$ and $\phi_{0}$ are sampled from flat distributions.
@@
As the next step, we estimate the confusion noise produced by these gravitational-wave sources in our mock Milky Way using the pipeline [^2] https://gitlab.in2p3.fr/Nikos/gwg. described in Karnesis et al. (2021, see also ). The pipeline approximates the so-called “global fit” analysis, which is the currently adopted approach to handling LISA’s complex data analysis (Colpi et al., 2024, see also Littenberg & Cornish 2023; Katz et al. 2024; Deng et al. 2025). The pipeline employs a signal-to-noise ($\rho$) evaluation to iteratively estimate the strain amplitude spectral density resulting from the combined signals of the unresolved (i.e., low $\rho$) part of the input population. As a result, we also obtain a catalog of individually resolved (i.e., high $\rho$) binaries. For our analysis, we adopt LISA’s instrumental noise requirements as defined in the technical note by LISA Science Study Team (2018). We assume a mission duration of 4 yr and use an $\rho$ threshold of 7 to distinguish between individually resolvable LISA sources and unresolved ones, a choice commonly adopted in detailed simulations (e.g., Crowder & Cornish, 2007; Cornish & Littenberg, 2007; Finch et al., 2023)
@@
We estimate a confusion foreground from our mock population containing double white dwarf binaries from both isolated binary and triple channels to be at the level of the LISA instrument noise. We also obtain $\sim 1.7\times 10^{4}$ sources above $\rho$ threshold, of which $\sim 6.5\times 10^{3}$ come from the isolated binaries and $\sim 1.1\times 10^{4}$ come from the triples. We illustrate these in Figure 9 (see Section 5.1 for comparison to previous works). Our estimate of the confusion noise is lower than those presented in the LISA mission proposal (Amaro-Seoane et al., 2017) and the more recent LISA Definition Study Report (Colpi et al., 2024), which used a different population synthesis study (Nelemans et al., 2004; Korol et al., 2017). This discrepancy arises from multiple aspects of population modeling, including variations in binary evolution prescriptions, Milky Way modeling, and, importantly, the inclusion of the triple formation channel in our simulations. However, since the same systematics are applied to both our triple and isolated binary populations, their differential properties remain largely unaffected. These factors collectively influence the total number of binaries in the LISA band and their properties. Identifying a single source of the difference is challenging, as these aspects are interrelated and non-trivially correlated. The LISA Consortium’s Astrophysics Working Group is currently investigating the differences and uncertainties in predicting the confusion foreground as part of the Ultra-Compact Binaries catalog comparison project (Valli et al., 2023, as well as Breivik et al. and Bobrick et al. in prep.). We refer the reader to those forthcoming results and provide a comparison using our test simulation, in which all LISA binaries were generated solely via the isolated binary channel.
@@ 4.1.2 Eccentric systems
Here, we focus on eccentric systems originating from the triple formation channel. These systems were excluded from the analysis above as they are estimated to be very few (and therefore do not contribute to the overall Galactic confusion signal) and because their gravitational-wave signals differ from those of circular systems.
@@
In our simulations, all eccentric systems show high eccentricities ($e>0.9$) and wide semi-major axis ($10^{1}-10^{6}\,\rm au$). For such orbital configurations, gravitational-wave emission predominantly occurs near pericenter passage, lasting up to a few hours and producing burst-like gravitational-wave signals. Since the orbital periods of these systems in our simulations are significantly longer than the mission duration (¿ 46 yr), the double white dwarf burst signals will not repeat within LISA’s observation window of 4 yr.
The probability to observe a gravitational-wave burst from an individual system at periapsis is $\sim\mathrm{T_{obs}/T_{orb}}$, where $\mathrm{T_{obs}}$ is the observational time ( $\mathrm{\sim 4\,yr}$). Assuming Poisson binomial distribution, we estimate that LISA will detect at most one eccentric double white dwarf during its operational duration. Following Xuan et al. (2024), we estimate the frequency and strain amplitude of such a burst signal as
@@
where $r_{\rm p}$ is the periapsis distance. We find the (dimensionless) strain amplitudes of the gravitational-wave emitted during the periapsis are between $\sim 10^{-24}$ and $10^{-22}$, i.e. below the noise curve (black stars in Figure 9). Therefore, such signal would not be detectable by LISA.
Figure 9: Characteristic strain $h_{c}={\cal A}_{\rm gw}\sqrt{f_{\rm gw}T_{\rm obs}}$ of resolved double white dwarf binaries in our mock simulation for a mission duration of $T_{\rm obs}=4\,yr$: triples in green points and isolated binary channel in purple points. This is compared to the LISA instrumental noise (black solid line). Triples with highly eccentric orbits ($e>0.9$) and other low signal-to-noise sources are marked in black and dark gray markers, respectively. The confusion background from the Galactic double white dwarf population is represented by green, purple, and orange lines for triples, binaries, and their combined contribution, respectively. All the eccentric systems (black markers) tend to have narrow periapsis time ($10^{4}\,\mathrm{s}$) and hence they occupy a narrow frequency range. The astrophysical noise from these backgrounds remains significantly lower than the instrumental noise.

### v2

@@ 3.3 Ejected tertiary
About $27^{+4}_{-2}\,\%$ of Galactic double white dwarfs from triples lost the tertiary star. The unbinding of the tertiary companion primarily occurs due to the following reasons. First, during a CE phase the inner binary’s orbit shrinks and loses angular momentum, while the prompt mass loss during the CE unbinds the outer orbit. Second, if the inner binary widens due to mass transfer or stellar winds the system becomes less hierarchical and dynamically unstable. Dynamically unstable orbits lead to chaotic evolution which may eject an object from the system. We provide a specific example of this process below.
Panel (b) of Fig. 6 shows an example of a system where a tertiary gets ejected after a CE in the inner binary. The system starts with $m_{1}=1.39\,\mathrm{\mathrm{M_{\odot}}}$ and $m_{2}=1.39\,\mathrm{\mathrm{M_{\odot}}}$ as the inner binary components and $m_{3}=0.31\,\mathrm{\mathrm{M_{\odot}}}$ as the tertiary star. The inner binary is initially eccentric with an inner eccentricity $\mathrm{e_{1}}=0.56$ and an inner semi-major axis $\mathrm{a_{1}}=50.1\,\mathrm{au}$. The tertiary orbit is wide, with an outer semi-major axis $\mathrm{a_{2}}\approx 1690\,\mathrm{au}$ and is eccentric with $\mathrm{e_{2}}=0.81$. The mutual inclination between the inner and outer orbits is 112. The inner binary masses are greater than the tertiary star so they evolve on a shorter timescale. The CE occurs in the inner binary during the AGB phase after about 3.6 Gyr. The CE reduces the inner binary semi-major axis to $10^{-2}\,\rm{au}$ and unbinds the outer tertiary. The remaining binary components evolve into a double white dwarf system. This double white dwarf binary emits gravitational waves and later enters the LISA frequency bandwidth after $1\,\rm{Myr}$. In this example, even though the binary does not have a bound tertiary star by the time it enters the LISA frequency bandwidth, the tertiary plays a role in shrinking the binary before the CE episode.
Figure 7: Schematic diagram for an example system that undergoes a TCE. In this scenario, the massive star transfers mass onto the inner binary, leading to its merger and the formation of a rejuvenated star. This rejuvenated star later enters the LISA frequency bandwidth along with the tertiary star. The legends are similar to those in Fig. 2. In addition, the yellow circle represents a star in the asymptotic giant branch phase.
@@ 3.4 Triple common envelope
About $5.5^{+2}_{-1}\,\%$ of Galactic double white dwarfs from triples undergo a phase of TCE evolution before entering the LISA frequency band. We identified two possible evolutionary outcomes of a TCE event in our simulations. First, the merger of the inner binary: the inner binary may merge to form a rejuvenated star. The new star forms a binary with the third star which later enters LISA frequency bandwidth. Second, ejection of one component: One of the stars may become unbound from the system, leaving behind a binary formed by the remaining two stars. This binary may enter the LISA band. We explain one of the examples below (see 2022ApJS..259...25H for a more detailed study of TCE).
Figure 7 shows an example of a TCE where the inner binary merges to form a new binary. The system starts with $m_{1}=1.18\,\mathrm{\mathrm{M_{\odot}}}$, $m_{2}=0.54\,\mathrm{\mathrm{M_{\odot}}}$, and a massive tertiary $m_{3}=5.59\,\mathrm{M_{\odot}}$. The inner binary is initially circular, with a semi-major axis $\mathrm{a_{1}}=0.05\,\mathrm{au}$. The tertiary orbit is compact, with an outer semi-major axis $\mathrm{a_{2}}=25\,\mathrm{au}$ and highly eccentric with $\mathrm{e_{2}}=0.94$. The mutual inclination between the inner and outer orbit is 160. Since the tertiary star is massive compared to the inner binary masses, it is the first to reach the AGB phase, in about 92 Myr. The inner binary components are still on the main-sequence.
@@
After approximately $2\,\rm{Gyr}$, the rejuvenated star evolves into a red giant which transfers mass onto the white dwarf, initiating another CE phase. The end of this CE phase leaves a circular and compact double white dwarf in a circular orbit with a semi-major axis of $\sim 8\times 10^{-2}\,\mathrm{au}$. After a few million years the binary enters the LISA frequency bandwidth due to gravitational wave emission. Here, TCE plays a significant role as it leads to merging the inner binary. The binary with the post-merger rejuvenated star later enters the LISA frequency bandwidth. Hence, the tertiary star plays a vital role in the evolution of the triple to LISA frequency bandwidth.
Table 1: Estimated counts.
Category $N(f<10^{-4}\,\textrm{Hz})\times 10^{6}$ $N(\rho>7)\times 10^{3}$ Triple $7.20^{+0.52}_{-0.50}$ $10.9^{+0.67}_{-0.50}$ Binary $3.80^{+0.17}_{-0.59}$ $6.5^{+0.22}_{-1.19}$ Triple + Binary $11.00^{+0.69}_{-1.09}$ $17.4^{+0.89}_{-1.69}$ Only binaries ($f_{t}=0$) $9.40^{+0.90}_{-0.67}$ $14.4^{+0.76}_{-0.45}$
 [^2] Table presents the estimated counts of total and individually resolvable (signal-to-noise ratio $\rho>7$) Galactic LISA double white dwarfs from different formation channels: Triple (originating from triple systems), Binary (originating from isolated binaries), Triple + Binary (the combined population from both channels), and Only binaries (isolated binaries assuming no contribution from triples, i.e., a triple fraction $f_{t}=0$). The quoted uncertainty estimates correspond to a 68.3% confidence interval and are estimated from bootstrap resampling and capture the statistical uncertainty introduced by the stochastic seeding of the Galaxy (see Appendix B for details).
@@ 3.5 Inner binary channel
We evolve the inner binaries of triple systems in isolation (i.e., without the influence of the third star) to quantify the impact of tertiary companions. About $39^{+8}_{-5}\,\%$ of the Galactic double white dwarfs from triples in our simulations enter LISA frequency bandwidth without any notable contribution from the third star. However, the initial configurations of these binaries (particularly their shorter orbital periods) are themselves a consequence of the dynamical stability constraints imposed by the tertiary companion. Here, the inner binary undergoes phases of CE before forming a short-period double white dwarf. This double white dwarf later enters the LISA frequency bandwidth either with a wide tertiary or as a binary with an ejected tertiary (for details on ejected tertiary, see Sect. 3.3).
4 LISA double white dwarfs: Isolated binary versus triple evolution
Building on the method in Sect. 2.3, where we construct the Galactic population by combining MSE with a Milky Way-like galaxy from TNG50, we obtain a total of $\sim 7.2\times 10^{6}$ Galactic double white dwarfs originating from triples. To compare this result with the isolated binary channel, we adjust our simulation as described in Appendix A. We emphasize that the initial distribution of the isolated binary population is constructed to represent truly isolated binaries and is different from the initial properties sampled to assemble the inner binaries in triple systems (See Sect. 2.2). In the isolated binary case, we find that about $\sim 3.8\times 10^{6}$ double white dwarfs are produced through isolated binary evolution, yielding a total of $\sim 1.1\times 10^{7}$ Galactic double white dwarfs currently emitting gravitational waves within the LISA bandwidth. Thus, approximately $65\,\%$ ($\sim 7.2\times 10^{6}$) of all Galactic double white dwarf binaries originated from triple systems as illustrated in Fig. 8. We note that the isolated binary evolution channel only yields circular binaries. In contrast, the triple channel generates a small fraction ($3\times 10^{-6}$) of eccentric binaries, which we discuss below.
Notably, we find that among all the double white dwarfs initially in triples only about half ($\sim$57%) retain a bound tertiary. In the remaining systems, the third star either became unbound or there was a merger of inner binary stars, reducing the system from a triple to a binary (See Sects. 3.3 and 3.2). Importantly, all double white dwarfs that retain their tertiary companion in our simulations have the tertiary in a wide orbit. We discuss the potential for detecting the presence of the tertiary companion based on LISA data in Sect. 4.4. We recall that we also generated a comparative simulation where the Galactic population only follows the isolated binary channel (i.e., entirely excluding triple channels). In this simulation we obtain a total of $\sim 9.4\times 10^{6}$ double white dwarfs emitting at LISA frequencies. All results are summarized in Table 1.
@@ 4.1.1 Circular systems
Double white dwarfs in the LISA band are typically millions of years away from merging. They are continuous, quasi-monochromatic gravitational wave sources for LISA. Describing these signals requires a set of eight parameters, typically chosen as $\{{\mathcal{A}_{\rm gw},f_{\rm gw},\dot{f}_{\rm gw},\lambda,\beta,\iota,\psi,\phi_{0}}\}$ (LISAWaveform). Here, $\mathcal{A}_{\rm gw}$ represents the gravitational wave amplitude, $f_{\rm gw}$ and $\dot{f}_{\rm gw}$ denote the gravitational wave frequency and its time derivative (or chirp), $(\lambda,\beta)$ correspond to the ecliptic longitude and latitude, respectively, $\iota$ is the (inner) binary inclination angle with respect to the line-of-sight, $\psi$ is the gravitational wave polarization angle, and $\phi_{0}$ represents the binary’s initial phase. Our population synthesis models provide binary parameters such as component masses, orbital periods and eccentricities, sky positions, and distances. We use these to derive gravitational wave parameters as follows. As discussed above, the overwhelming majority of double white dwarfs in our simulations are circularized.
The gravitational wave frequency of a circular binary is twice its orbital frequency ($f_{\rm orb}$),
@@
where $G$ and $c$ are the gravitational constant and speed of light respectively. The amplitude is set by the source’s distance, $d$, and chirp mass:
@@
Equations (3), (4), and (6) define the first three parameters of the set. The ecliptic coordinates $(\lambda,\beta)$ are inherited from the TNG50 Milky Way-like galaxy where the binary was seeded, while the remaining three parameters are assigned randomly. Specifically, $\iota$ is sampled from a uniform distribution in $\cos\iota$, and $\psi$ and $\phi_{0}$ are sampled from flat distributions.
As the next step, we estimate the confusion noise produced by these gravitational wave sources in our mock Milky Way using the pipeline [^3] https://gitlab.in2p3.fr/Nikos/gwg. described in kar21. The pipeline approximates the so-called “global fit” analysis, which is the currently adopted approach to handling LISA’s complex data analysis (LISARedBook, see also Littenberg_2023; katz2024efficientgpuacceleratedmultisourceglobal; deng2025modularglobalfitpipelinelisa).
@@
The pipeline employs a signal-to-noise ($\rho$) evaluation to iteratively estimate the strain amplitude spectral density resulting from the combined signals of the unresolved (i.e., low $\rho$) part of the input population. As a result, we also obtain a catalog of individually resolved (i.e., high $\rho$) binaries. For our analysis, we adopt LISA’s instrumental noise requirements as defined in the technical note by LISAdoc. We assume a mission duration of 4 yr and use an $\rho$ threshold of 7 to distinguish between individually resolvable LISA sources and unresolved ones, a choice commonly adopted in detailed simulations (e.g., cro07; Cornish_2007; 2023MNRAS.522.5358F)
We estimate a confusion foreground from our mock population containing double white dwarf binaries from both isolated binary and triple channels to be at the level of the LISA instrument noise. We also obtain $\sim 1.7\times 10^{4}$ sources above the $\rho$ threshold, of which $\sim 6.5\times 10^{3}$ come from the isolated binaries and $\sim 1.1\times 10^{4}$ come from the triples (see Sect. 5.1 for comparison to previous works). Figure 9 shows the characteristic strain as a function of gravitational wave frequency for Galactic double white dwarf sources originating from both triples and isolated binaries. The characteristic strain is defined as the effective strain amplitude, incorporating the number of cycles a system completes during the observation time. This quantity is convenient for comparing source signals with the detector sensitivity, since the height of a system above the sensitivity curve represents its signal-to-noise ratio.
Our estimate of the confusion noise is lower than that presented in the LISA mission proposal (LISAproposal) and the more recent LISA Definition Study Report (LISARedBook), which used a different population synthesis study (2004MNRAS.349..181N; 2017MNRAS.470.1894K). This discrepancy arises from multiple aspects of population modeling, including variations in binary evolution prescriptions, Milky Way modeling, and, importantly, the inclusion of the triple formation channel in our simulations. However, since the same systematics are applied to both our triple and isolated binary populations, their differential properties remain largely unaffected.
@@
These factors collectively influence the total number of binaries in the LISA band and their properties. Identifying a single source of the difference is challenging, as these aspects are interrelated and nontrivially correlated. The LISA Consortium’s Astrophysics Working Group is currently investigating the differences and uncertainties in predicting the confusion foreground as part of the Ultra-Compact Binaries catalog comparison project (2023arXiv231103431V, as well as Breivik et al. and Bobrick et al. in prep.). We refer the reader to those forthcoming results and provide a comparison using our test simulation, in which all LISA binaries were generated solely via the isolated binary channel.
@@ 4.1.2 Eccentric systems
Here, we focus on eccentric systems originating from the triple formation channel. These systems were excluded from the analysis above as they are estimated to be very few (and therefore do not contribute to the overall Galactic confusion signal) and because their gravitational wave signals differ from those of circular systems.
@@
In our simulations, all eccentric systems show high eccentricities ($e>0.9$) and wide semi-major axis ($10^{1}-10^{6}\,\rm au$). For such orbital configurations, gravitational wave emission predominantly occurs near pericenter passage, lasting up to a few hours and producing burst-like gravitational wave signals. Since the orbital periods of these systems in our simulations are significantly longer than the mission duration (¿ 46 yr), the double white dwarf burst signals will not repeat within LISA’s observation window of 4 yr.
The probability to observe a gravitational wave burst from an individual system at periapsis is $\sim\mathrm{T_{obs}/T_{orb}}$, where $\mathrm{T_{obs}}$ is the observational time ( $\mathrm{\sim 4\,yr}$) and $\mathrm{T_{orb}}$ is the orbital period of the binary. Assuming Poisson binomial distribution, we estimate that LISA will detect at most one eccentric double white dwarf during its operational duration. Following 2024ApJ...965..148X, we estimated the frequency and strain amplitude of such a burst signal as
@@
where $r_{\rm p}$ is the periapsis distance. We found the (dimensionless) strain amplitudes of the gravitational wave emitted during the periapsis are between $\sim 10^{-24}$ and $10^{-22}$, i.e., below the noise curve (black stars in Fig. 9). Therefore, such a signal would not be detectable by LISA.
Figure 9: Characteristic strain $h_{c}={\cal A}_{\rm gw}\sqrt{f_{\rm gw}T_{\rm obs}}$ of resolved double white dwarf binaries in our mock simulation for a mission duration of $T_{\rm obs}=4\,yr$. Triples are shown with green points, and the isolated binary channel is shown with purple points. This is compared to the LISA instrumental noise (black solid line). Triples with highly eccentric orbits ($e>0.9$) and other low signal-to-noise sources are marked in black and dark gray markers, respectively. The confusion background from the Galactic double white dwarf population is represented by green, purple, and orange lines for triples, binaries, and their combined contribution, respectively. All the eccentric systems (black markers) tend to have a narrow periapsis time ($10^{4}\,\mathrm{s}$), and hence they occupy a narrow frequency range. The astrophysical noise from these backgrounds remains significantly lower than the instrumental noise.